Daddi Mac, Inc. doesn’t face any taxes and has $313.00 million in assets, currently financed entirely with equity. Equity is worth $40 per share, and book value of equity is equal to market value of equity. Also, let’s assume that the firm’s expected values for EBIT depend upon which state of the economy occurs this year, with the possible values of EBIT and their associated probabilities as shown below:

State Recession Average Boom
Probability of state 0.25 0.55 0.20
Expected EBIT in state $ 6,134,800 $ 11,080,200 $ 18,091,400

  

The firm is considering switching to a 20-percent-debt capital structure and has determined that it would have to pay an 8 percent yield on perpetual debt in either event. What will be the level of expected EPS if the firm switches to the proposed capital structure?
Answer #1

We have to find EPS for each of the state given
Interest expense=debt*interest rate
=313*10^6*20%*8%=5008000
Net income=EBIT-interest expenses
Recession :=6134000-5008000=1126800
Average:=11080200-5008000=6072200
Boom:=18091400-5008000=13083400

New no of shares after change in capital structure= equity value/market price
=(313*10^6*80%)/40=6260000
EPS in each state below= Net income/shares outstanding
Recession:=1126800/6260000=0.18
Average:=6072200/6260000=0.97
Boom:=13083400/6260000=2.09

Expected EPS=sum of (prob*EPS) in each state
=(0.25*0.18)+(0.55*0.97)+(0.2*2.09)=0.9965
rounded to 2 decimals it is 1.00